OpenStudy (anonymous):

Evaluate \(\frac{d}{d x} \int_{t = 0}^{t = x^2} \cos (t^2) d t =\)?\(\cos (\)?\()\)

OpenStudy (zarkon):

\[\cos((x^2)^2)2x=2x\cos(x^4)\]